# Unstable entropy for Anosov diffeomorphisms on the $3$-torus

### v1

@@ 4.3. Dynamical determinants
where we used (1.6) for $\lambda=\lambda_{u}$ and $u=\theta$. The rest of the argument follows the proof of [12, Lemma 8.1]. We will need the following characterization of the conditional measures, see [12, Proposition 8.2].
@@
Let $x\in\mathbb{T}^{3}$, $n\geq 0$ and $R\subset\mathbb{T}^{3}$ be a rectangle so that $\mu_{J^{c}_{f}}(R)>0$. Let $\{\xi_{n}\}_{n\in\mathbb{N}}$ be a sequence of refining finite partitions so that for any $y\in R$, $\xi_{n}(y)$ is a rectangle and $\cap_{n\in\mathbb{N}}\xi_{n}(y)=\mathcal{W}^{cs}_{R}(y)$. Note that $(f^{*})^{n}\theta_{x}$ is supported on $\mathcal{W}^{cs}(f^{-n}x)$. Fix $y\in R^{\prime}$ given by Lemma 5.4 and let $n,m\in\mathbb{N}$. Then there exists points $z^{(1)}_{n,m},\ldots,z^{(s(m))}_{n,m}\in\mathcal{W}^{cs}(f^{-n}x)\cap\xi_{m}(y)$ such that $\mathcal{W}^{cs}(f^{-n}x)\cap\xi_{m}=\cup_{i=1}^{s(m)}\mathcal{W}^{cs}_{R}(z^{(i)}_{n,m})$, where the union is disjoint [^4] Since our system $(\theta_{x})_{x\in\mathbb{T}^{3}}$ is defined on the whole center stable manifold $\mathcal{W}^{cs}(x)$ rather than the local one, any $\mathcal{W}^{cs}_{R}(z^{(i)}_{n,m})$ is fully included in $\mathcal{W}^{cs}(f^{-n}x)$ in contrast to [12, Proposition 8.2]. . We want to write $((f^{*})^{n}\theta_{x})_{|\xi_{m}}$ as a linear combination of measures supported on these sets. Using (5.2) gives, for any continuous $\psi:\xi_{m}(y)\to\mathbb{R}$,
@@
We are left with showing the wavefront set bound. The proof follows [27, Lemma 3.2]. Let $\chi$ be a smooth $2$-form supported in $R$ and $S\in C^{\infty}$ such that $dS(q)=\xi\notin E_{cs}^{*}.$
@@
As explained in Remark 1, the condition $\mathrm{WF}({\mu^{u}})\subset E_{cs}^{*}$ is stronger than the condition in (1.7). Lemma 4.6 and [29, Lemma 4.3.2] imply that that $\mathrm{WF}(\mu_{J^{c}_{f}})\subset E_{cs}^{*}\oplus E_{cu}^{*}$. This is better than the usual wavefront set bound $\mathrm{WF}(\mu)\subset E_{s}^{*}\oplus E_{cu}^{*}$ for an equilibrium state $\mu$. This extra regularity in the center direction is actually a consequence of the fact that the conditional densities of $\mu_{J^{c}_{f}}$ along the center direction are absolutely continuous with respect to the Lebesgue measure. Indeed, by Proposition 4.1, one has equality in (2.11) and thus in (2.12) and [38, Theorem 3.7] gives that the conditional densities of $\mu_{J^{c}_{f}}$ on center manifolds are absolutely continuous. This should be compared to the case of the SRB measure for which $\mathrm{WF}(\mu_{\mathrm{SRB}})\subset E_{cu}^{*}$ since its conditional densities on the stable manifolds are absolutely continuous.

### v3

@@ 4.3. Dynamical determinants
where we used (1.6) for $\lambda=\lambda_{u}$ and $u=\theta$. The rest of the argument follows the proof of [undefk, Lemma 8.1]. We will need the following characterization of the conditional measures, see [undefk, Proposition 8.2].
@@
Let $x\in\mathbb{T}^{3}$, $n\geq 0$ and $R\subset\mathbb{T}^{3}$ be a rectangle so that $\mu_{J^{c}_{f}}(R)>0$. Let $\{\xi_{n}\}_{n\in\mathbb{N}}$ be a sequence of refining finite partitions so that for any $y\in R$, $\xi_{n}(y)$ is a rectangle and $\cap_{n\in\mathbb{N}}\xi_{n}(y)=\mathcal{W}^{cs}_{R}(y)$. Note that $(f^{*})^{n}\theta_{x}$ is supported on $\mathcal{W}^{cs}(f^{-n}x)$. Fix $y\in R^{\prime}$ given by Lemma 5.4 and let $n,m\in\mathbb{N}$. Then there exists points $z^{(1)}_{n,m},\ldots,z^{(s(m))}_{n,m}\in\mathcal{W}^{cs}(f^{-n}x)\cap\xi_{m}(y)$ such that $\mathcal{W}^{cs}(f^{-n}x)\cap\xi_{m}=\cup_{i=1}^{s(m)}\mathcal{W}^{cs}_{R}(z^{(i)}_{n,m})$, where the union is disjoint [^4] Since our system $(\theta_{x})_{x\in\mathbb{T}^{3}}$ is defined on the whole center stable manifold $\mathcal{W}^{cs}(x)$ rather than the local one, any $\mathcal{W}^{cs}_{R}(z^{(i)}_{n,m})$ is fully included in $\mathcal{W}^{cs}(f^{-n}x)$ in contrast to [undefk, Proposition 8.2]. . We want to write $((f^{*})^{n}\theta_{x})_{|\xi_{m}}$ as a linear combination of measures supported on these sets. Using (5.2) gives, for any continuous $\psi:\xi_{m}(y)\to\mathbb{R}$,
@@
We are left with showing the wavefront set bound. The proof follows [undefz, Lemma 3.2]. Let $\chi$ be a smooth $2$-form supported in $R$ and $S\in C^{\infty}$ such that $dS(q)=\xi\notin E_{cs}^{*}.$
@@
As explained in Remark 1, the condition $\mathrm{WF}({\mu^{u}})\subset E_{cs}^{*}$ is stronger than the condition in (1.7). Lemma 4.6 and [undefab, Lemma 4.3.2] imply that that $\mathrm{WF}(\mu_{J^{c}_{f}})\subset E_{cs}^{*}\oplus E_{cu}^{*}$. This is better than the usual wavefront set bound $\mathrm{WF}(\mu)\subset E_{s}^{*}\oplus E_{cu}^{*}$ for an equilibrium state $\mu$. This extra regularity in the center direction is actually a consequence of the fact that the conditional densities of $\mu_{J^{c}_{f}}$ along the center direction are absolutely continuous with respect to the Lebesgue measure. This should be compared to the case of the SRB measure for which $\mathrm{WF}(\mu_{\mathrm{SRB}})\subset E_{cu}^{*}$ since its conditional densities on the stable manifolds are absolutely continuous.
Proposition 5.6.
The conditionals of $\mu_{J^{c}_{f}}$ on the center foliation $(\mathcal{W}^{c}(x))_{x\in\mathbb{T}^{3}}$ are absolutely continuous with respect to the Lebesgue measures $d\mathrm{Vol}_{\mathcal{W}^{c}(x)}$.
Proof.
For $z\in\mathbb{T}^{3}$ and $y\in\mathcal{W}^{c}(z)$, we define
(5.14) $\omega(z,y):=\sum_{k=0}^{+\infty}\big(J^{c}_{f}(f^{-k}y)-J^{c}_{f}(f^{-k}z)\big),$
where the sum converges since $\mathcal{W}^{c}$ is uniformly expanding and $J^{c}_{f}$ is Hölder continuous. We first use that the strong unstable foliation is $C^{1}$ when restricted to the unstable foliation, see for instance [undef, Section 2.2.1]. In particular, that there is $\delta>0$, such that for any $x\in\mathbb{T}^{3}$ and $y\in\mathcal{W}^{u}(x,\delta)$, one has $(\mathrm{Hol}^{u}_{x,y})_{*}\mathrm{Vol}_{\mathcal{W}^{c}(x)}\ll\mathrm{Vol}_{\mathcal{W}^{c}(y)}$. The density can be computed using the fact that $d(f^{-k}x,f^{-k}y)\to 0$ when $k\to+\infty$ and the fact that the holonomy is $C^{1}$. A direct computation using Jacobians shows that for any $z\in\mathbb{T}^{3}$ and $y\in\mathcal{W}^{c}(z)$, the density is given by the exponential of the function $\omega$ defined in (5.14)
(5.15) $\forall u\in\mathcal{W}^{c}(y),\quad\frac{d\big((\mathrm{Hol}^{u}_{z,y})_{*}\mathrm{Vol}_{\mathcal{W}^{c}(z)}\big)}{d\mathrm{Vol}_{\mathcal{W}^{c}(y)}}(u)=e^{\omega(u,\mathrm{Hol}^{s}_{y,z}u)}.$
We now define a system of measures on the unstable foliation $\mathcal{W}^{cu}$. For any $x\in\mathbb{T}^{3}$ and any compactly supported $\phi\in C(\mathcal{W}^{cu}(x))$, we let
(5.16) $\mu^{cu}_{x}(\phi)=\int_{\mathcal{W}^{u}(x)}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(z,y)}\phi(z)d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)d\mu^{u}_{x}(y).$
The definition does not depend on the choice of base point $x$ in the unstable leaf by (5.15) and the fact that $\mu^{u}$ does not depend on the choice of base point in the strong unstable leaf. Since $\mu^{u}$ was shown to be a Radon measure with full support on each strong unstable leaf, see Theorem 2, this shows that $\mu^{cu}$ is also a Radon measure with full support on each unstable leaf. We now compute, using $J^{c}_{f}(z)=\omega(f(z),z)$ and the cocycle relation $\omega(x,y)+\omega(y,z)=\omega(x,z)$,
$\mu^{cu}_{x}(e^{-J^{c}_{f}(\cdot)}\phi\circ f)$ $=\int_{\mathcal{W}^{u}(x)}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(z,y)}e^{-J^{c}_{f}(z)}\phi(f(z))d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)d\mu^{u}_{x}(y)$
$=\int_{\mathcal{W}^{u}(x)}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(z,y)}e^{\omega(f(z),z)}\phi(f(z))d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)d\mu^{u}_{x}(y)$
$=\int_{\mathcal{W}^{u}(x)}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(f(z),y)}\phi(f(z))d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)d\mu^{u}_{x}(y)$
$=\int_{\mathcal{W}^{u}(x)}f^{*}\big(\mathrm{Vol}_{\mathcal{W}^{c}(y)}(e^{\omega(\cdot,y)}\phi(\cdot))\big)d\mu^{u}_{x}(y)$
$=\int_{\mathcal{W}^{u}(f(x))}\left(\int_{\mathcal{W}^{c}(y)}e^{\omega(z,y)}e^{-J^{c}_{f}(z)}\phi(f(z))d\mathrm{Vol}_{\mathcal{W}^{c}(y)}(z)\right)df_{*}(\mu^{u}_{x})(y)$
$=e^{-h^{u}_{\mathrm{top}}(f)}\mu_{f(x)}^{cu}(\phi)=e^{-P(J^{c}_{f})}\mu_{f(x)}^{cu}(\phi),$
where we used that $h^{u}_{\mathrm{top}}(f)=P(J^{c}_{f})$ by Theorem 1 and the conformal property of $\mu^{u}$, see Theorem 2.
In other words, we checked that the system $(\mu^{cu}_{x})_{x\in\mathbb{T}^{3}}$ is a $cu$-system which satisfies properties $(2)$ and $(3)$ of [undefi, Theorem A] for the potential $J^{c}_{f}$. We can thus apply [undefi, Theorem B] which gives that the $cu$-conditionals of the equilibrium state $\mu_{J^{c}_{f}}$ associated to $J^{c}_{f}$ are absolutely continuous with respect to $\mu^{cu}$.
From (5.15), the $c$-conditionals of $\mu^{cu}$ are absolutely continuous with respect to the Lebesgue measure on $\mathcal{W}^{c}$ and we deduce that this is also the case for $\mu_{J^{c}_{f}}$. ∎